OpenStudy (joclyn16):

A triangular prism is 16.6 inches long. It has a triangular face with a heightof 18 inches. The volume of the prism is 2554.74 in.³. What is the base of its triangular face?

OpenStudy (joclyn16):

Hurry

OpenStudy (anonymous):

Start with the volume formula for a prism.

OpenStudy (joclyn16):

So 16.6 times 18?

OpenStudy (anonymous):

V = 1/2 Base*Height*Length correct?

OpenStudy (anonymous):

Then fill in the things you know.

OpenStudy (joclyn16):

Ok

OpenStudy (joclyn16):

I just need the answer

OpenStudy (joclyn16):

Tell me the answer

OpenStudy (anonymous):

Here is a start 2554.74 = Base*18*Length

OpenStudy (joclyn16):

So what do I do

OpenStudy (anonymous):

Do you know the Length?

OpenStudy (joclyn16):

Yes 16.6

OpenStudy (anonymous):

Ok, then 2554.74 = Base*18*16.6

OpenStudy (joclyn16):

18 times 16.6 equals 298.8